# JoIN: Joint GANs Inversion for Intrinsic Image Decomposition

### v2

@@ 3.2 Intrinsic image decomposition
In this paper we focus on the problem of intrinsic image decomposition. We thus adapt the general formulation presented in the previous section.
We first consider a simple forward model assuming Lambertian surfaces. It uses two components – albedo and shading to represent the image:
$I^{*}=sRGB(C_{albedo}*C_{shading}).$ (5)
@@
$I^{*}=sRGB(C_{albedo}*C_{shading}+C_{specular}).$ (6)
@@ 3.3 Optimization-based Joint Inversion
Given a forward model described in Eq. 5 or Eq. 6, we first train a GAN for each intrinsic component (albedo, shading and eventually specular images) independently using the StyleGAN2 architecture [36, 38]. As the GANs are trained independently the forward model can be chosen after the fact and components replaced or added, we also demonstrate in Sec. 5.4 that training independent GANs helps prevent channel cross-contamination.
@@
Allowing the latent codes $w_{i}$ to explore freely the space during optimization can lead to cross-contamination as they deviate from the manifold of possible $w$ within $\mathcal{W}$. In such case the generator for a component $C_{i}$ starts producing signals corresponding to another component $C_{j}$ We thus propose a new kNN-based regularization loss to constrain the latent codes to stay within domain.
To this end, we take inspiration from previous works [87, 71], which introduces an in-domain loss on the latent codes as follows:
@@
where $\widebar{w}$ is the average $w$ obtained by drawing $100k$ random samples of random normal code $z$. This in-domain loss tries to keep the optimized latent close to the distribution mean in order to keep the estimate within the training domain. We argue that this formulation has two main drawbacks. First, using the distribution mean penalizes the solution space near the boundary of the distribution. Second, it assumes that the distribution of $w$ is isotropic, which has no reason to hold given the use of a non-linear mapping network during training. As a remedy, we propose a novel k-Nearest Neighbors-based loss formulation that better allows the exploration of the full domain. We first sample a large number ($100k$) of $z$ codes from a truncated Gaussian distribution to obtain a set of $w$ codes $W$.
@@
Then, during inversion, at each step, we calculate the $k$ nearest neighbors of the current latent code estimate $kNN(w_{i})$, and use a softmax weighted average of the distance to nearest neighbors as a loss:
$\mathcal{L}_{kNN}=\sum_{w_{i}^{k}\in kNN(w_{i})}\|w_{i}-w_{i}^{k}\|_{2}\cdot\frac{e^{-0.5\frac{\|w_{i}-w_{i}^{k}\|_{2}}{\widebar{d}^{k}_{i}}}}{D_{i}}$ (8)
where $D_{i}$ is the normalizing factor:
$D_{i}=\sum_{w_{i}^{k}\in kNN(w_{i})}e^{-0.5\frac{\|w_{i}-w_{i}^{k}\|_{2}}{\widebar{d}^{k}_{i}}}$ (9)
@@
And $\widebar{d}^{k}_{i}$ is the mean distance to nearest neighbors:
$\widebar{d}^{k}_{i}=\sum_{w_{i}^{k}\in kNN(w_{i})}\|w_{i}-w_{i}^{k}\|_{2}$ (10)
@@
Figure  2 provides an intuitive visualization of our kNN loss in 2D. Our loss landscape does not penalize as much the exploration of the distribution and its boundaries. We provide the results of our optimization scheme and ablation studies demonstrating the benefits of our loss in Sec. 5.
Figure 2: Loss landscapes for 2D latent vectors. Dots represent randomly sampled latent vectors and the colormap the intensity of the loss. 100 samples, k=5 NN. Left: using in-domain formulation. Right: using our knn loss. Ours allows to better capture the diversity of the distribution and is lower at boundaries.
@@ 4 Adapting existing inversion methods
In this section we show how existing initialization and generator fine-tuning mechanisms can be adapted to JoIN to improve the inversion result.
@@ 4.1 Encoder-guided Initialization
While our optimization-based approach produces faithful decompositions on synthetic data, the optimization can get stuck in local minima if the initial latent codes lie far from the target codes. To make our approach more robust, we propose to use an encoder-guided initialization inspired by hybrid approaches of GAN inversion [16, 8, 73]. For each component, we can independently train a pSp network  [63] to encode a natural image into a latent code $w$ used by the pre-trained generator to recover the given component. At test time, given a natural image, the latent code produced by the encoder can be used as an initialization for the optimization.
4.2 Generator fine-tuning - PTI
Recently, approaches using generator fine-tuning have emerged [65, 25]. First, a $w-$code is estimated using an off-the-shelf inversion technique, and then with the fixed $w$ pivot, the generator is updated to obtain close to zero reconstruction errors.
In our case, the generator fine-tuning scheme can be defined as:
$\operatorname*{arg\,min}_{\theta_{G_{i}},i\in{1\dots n}}\mathcal{L}\left(I^{*},f\left(T^{-1}_{1}\left(G_{1}\left(\widehat{w_{1}}\right)\right),\dots,T^{-1}_{n}\left(G_{n}\left(\widehat{w_{n}}\right)\right)\right)\right)$ (11)
@@ 4.2 Generator fine-tuning - PTI
While direct fine-tuning allows the generator to overfit to the target image perfectly, it distorts the GAN manifold. In our case, allowing multiple GANs to fine-tune at once can result in loss of the priors and cross-contamination (see Fig. 13 w/o D Loss). Therefore, we rely on the discriminator loss to regularize the fine-tuning [25]. While fine-tuning the generator around the $\widehat{w_{i}}$ pivot, we apply a new localized discriminator loss $\mathcal{L}_{LD}$, we refer to it as D Loss.
Specifically, we sample $n_{a}$ number of anchor latent codes in the close vicinity of our fixed $w-$code by interpolating $\widehat{w_{i}}$ with randomly sampled codes $w_{k}$. At each fine-tuning step, we pass the anchor codes through the generator and compute the discriminator loss on the resulting images, keeping the discriminator frozen:
@@
where $\mathcal{L}_{G}$ is the generator loss as defined in Eq. 11. As discussed in the Ablation study (Sec. 5.4), the $\mathcal{L}_{LD}$ loss helps maintaining the GAN priors during optimization even as multiple generators are fine-tuned together.
Figure 3: Decomposition results for synthetic Materials. Three examples are shown. Again, left column, from top to bottom, target Image (used for optimization), GT shading, and albedo. The right column contains our estimates. We are able to correctly separate the highlights and to discriminate between darker shading and darker albedo.
@@ 5 Experiments
In this section, we present a set of experiments that showcase the efficacy of our method. We provide intrinsic image decomposition results on two different datasets on both synthetic and real images along with comparisons with existing methods and ablation studies.
Figure 4: Comparisons for Albedo estimation on real material pictures. Our GAN priors allow us to better separate the highlight from the albedo leading to fewer visual artifacts. Our recovered specular layer is shown on the right for reference.
@@ 5.1 Preliminaries
For all our experiments we use StyleGAN2 [38] to train the generators and a pSp [64] encoder for initialization. All our generators and encoders are trained using their publicly available codebases and with the standard choices of the hyperparameters.
@@
The decomposition results are evaluated using several metrics such as MSE, SSIM, LPIPS, and PSNR for the synthetic images that provide ground truth. We also provide comparisons with several state-of-the-art methods that are used in the literature for intrinsic image decomposition. The results demonstrate the effectiveness of our proposed method in decomposing the input images into their intrinsic components and the ability to handle real images.
@@ 5.2 Experiments on Materials data
The Materials dataset is composed of highly realistic renderings of synthetic material tiles as done in Deschaintre et al. [19]. We obtain the albedo map directly from the material and compose the shading and specular components from blender direct and indirect rendering passes. More details on this dataset are given in the supplementary. For this dataset, we show results for decomposition into 3 components (albedo, shading, and specular; Eq. 6) and operate at the resolution of $512\times 512$. For synthetic test images, we show results in Fig. 3 and quantitative comparisons to several recent SVBRDF estimation methods [19, 84, 85] on the task of albedo recovery in Tab. 1. While we do not recover the full SVBRDF as these methods do, we do show competitive results on albedo map recovery both quantitatively and qualitatively on real images, taken from Deschaintre et al. [20], visible in Fig. 4. As can be seen, our GAN priors allow to correctly separate flash highlights from the diffuse component and to recover plausible content under the highlights in the albedo map while correcting for the lighting fall-off. Additional results are provided in the supplementary (Fig. 1617).
Table 1: Quantitative comparisons on synthetic Materials. We evaluate albedo estimation on 100 materials from our test dataset for our method and three recent SVBRDF estimation methods.
Estimated Albedo
MSEnormal-$\downarrow$ LPIPSnormal-$\downarrow$ PSNRnormal-$\uparrow$
Desch18 [19] 0.0211 0.314 19.53
Zhou21 [84] 0.0202 0.293 20.919
Zhou22 [85] 0.0213 0.296 20.18
Ours 0.0146 0.326 21.58
Figure 5: Two components decomposition results for Lumos Faces dataset. On synthetic data, our estimated components are close to the ground truth, generator tuning is not needed when testing data is close to the GAN training distribution.
@@ 5.3 Experiments on Faces data
Figure 6: Three components decomposition results for Lumos Faces dataset. Using three components allows to recover subtle specular effects.
Faces are notoriously adapted to the usage of GANs and to cater to such a use case, we use the synthetic Lumos dataset [79] to train our GANs, we defer more discussions on the dataset to the supplementary.
First, we show results on test images from the dataset in Fig. 5 and Fig. 6 where we use respectively two and three components. For this experiment, we use the encoder initialization and our optimization without generator fine-tuning (PTI). As can be seen in these figures our method is able to recover accurate albedo, shading, and specular maps and does not suffer from cross-channel contamination.
As the synthetic dataset does not reproduce the complexity and diversity of real faces, for real images, we add the generator fine-tuning step presented in Sec. 4.2. We show results on real faces taken from the FFHQ dataset in Fig. 7 and compare our method to five recent relighting methods, Total Relighting [57], Lumos [79], InverseRenderNet++ [81], PieNet [18], and Nextface [21, 22, 23], that seek to recover face albedo as part of their relighting pipeline, this is shown in Fig. 9. In all the cases, the predicted components by our method maintain the priors of each GAN. In the comparison figure we can observe that our method works better at producing shading-free albedo images. Notably, our predicted albedo contains fewer shading variations around the neck, nose, nostrils, and mouth.
For quantitative comparisons, we use a test set of synthetic images from Lumos dataset, and compare reconstruction errors for albedo recovery for all variants of our approach with four competing methods in Tab. 3. We see that our method achieves the lowest errors. We provide additional results on both synthetic and real faces in the supplementary (Fig. 18192021).
Figure 7: Decomposition results for real faces. Using generator fine-tuning our method generalizes to real faces and correctly separates albedo, shading, and specular effects while reconstructing the input correctly.
Figure 8: Face relighting results. Once the IID is obtained, we modify the shading component using GAN editing technique [68] on the shading generator obtaining plausible relighting edits.
Application to relighting. GANs are known for the rich image manipulation capabilities they offer. We show real image relighting results in Fig. 8. To generate these results, after obtaining the inversions for albedo, shading, and specular, we use an off-the-shelf latent space editing method SeFa [68] on the shading generator to modify the shading image while keeping the other two components constant. StyleGAN’s disentangled representation helps with identity preservation. As seen in Fig. 8, our approach produces varied novel lighting.
Figure 9: Comparison of our method on real face images. On real faces, our approach can recover more faithful albedo by removing shadows, and preserving the appropriate skin tones as compared to other competing methods.
Table 2: Quantitative comparisons between different variations of our approach on Lumos Faces synthetic dataset
Faces
Albedo Shading Specular Image
LPIPSnormal-$\downarrow$ MSEnormal-$\downarrow$ PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ MSEnormal-$\downarrow$ PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ MSEnormal-$\downarrow$ PSNRnormal-$\uparrow$ LPIPSnormal-$\downarrow$ MSEnormal-$\downarrow$ PSNRnormal-$\uparrow$
Inversion w/ single GAN for both albedo and shading 0.1954 0.1269 14.1662 0.2657 0.07021 18.2450 - - - 0.09342 0.03133 24.4665
Optimization 0.1311 0.0417 22.4763 0.2832 0.0518 18.8636 0.2544 0.0336 32.3207 0.0879 0.0308 24.2962
Opt. + InDomain loss 0.1209 0.0400 22.5891 0.2738 0.0497 19.1599 0.2444 0.0332 32.8339 0.0876 0.0302 24.1644
Opt. + kNN loss 0.1047 0.0357 23.5331 0.2603 0.0487 20.911 0.2361 0.0320 32.9050 0.0868 0.0301 24.1144
Encoder + Opt. + kNN loss 0.0887 0.0345 24.1467 0.2039 0.0414 21.0323 0.1722 0.0282 33.4090 0.0713 0.0306 26.4451
Encoder + Opt. + kNN loss + PTI w/o D Loss 0.0784 0.0155 24.1160 0.2775 0.0519 18.1651 0.1550 0.0278 32.8380 0.0136 0.0011 33.5454
Encoder + Opt. + kNN loss + PTI w/ D Loss 0.0765 0.0072 27.3401 0.1703 0.0285 22.4754 0.0774 0.0251 34.5475 0.0194 0.0022 32.5175
Table 3: Quantitative comparisons on Synthetic Faces testset
On synthetic faces (Lumos dataset) Estimated Albedo
LPIPSnormal-$\downarrow$ MSEnormal-$\downarrow$ PSNRnormal-$\uparrow$
Total Relighting [57] 0.2220 0.0644 17.1470
PieNet [18] 0.1843 0.0501 19.4761
InverseRenderNet++ [81] 0.3794 0.1550 14.5227
NextFace [21, 22, 23] 0.2034 0.0791 19.0699
@@
Opt. + kNN loss 0.1047 0.0357 23.5331
Encoder + Opt. + kNN loss 0.0887 0.0345 24.1467
Encoder + Opt. + kNN loss + PTI w/o D Loss 0.0784 0.0155 24.1160
Encoder + Opt. + kNN loss + PTI w/ D Loss 0.0765 0.0072 27.3401

### v3

@@ 3.2 Intrinsic Image Decomposition
In this paper, we focus on the problem of intrinsic image decomposition, covering several formulations (Garces et al., 2021). We thus adapt the general formulation presented in the previous section.
We first consider a simple forward model assuming Lambertian surfaces (Barrow & Tenenbaum, 1978). This seminal model uses two components – albedo and shading to represent the image:
$I^{*}=sRGB(C_{albedo}\cdot C_{shading}).$ (5)
@@
$I^{*}=sRGB(C_{albedo}\cdot C_{shading}+C_{specular}).$ (6)
This model merges the specular color and specular illumination of the dichromatic model Tominaga (1994) as a single component.
@@ 3.3 Optimization-based Joint Inversion
Given a forward model described in Eq. 5 or Eq. 6, we first train a GAN for each intrinsic component (albedo, shading and eventually specular images) independently using the StyleGAN2 architecture (Karras et al., 2019; 2020). As the GANs are trained independently the forward model can be chosen after the fact and components can be replaced or added.
@@
Allowing the latent codes $w_{i}$ to explore the latent space freely during optimization can lead to cross-contamination as they deviate from the manifold of possible $w$ within $\mathcal{W}$. In such case the generator for a component $C_{i}$ starts producing signals corresponding to another component $C_{j}$. We thus propose a new kNN-based regularization loss to constrain the latent codes to stay within domain.
To this end, we take inspiration from previous works (Zhu et al., 2020; Tov et al., 2021), which introduces an in-domain loss on the latent codes as follows:
@@
Figure 3: $\mathcal{W}$-space of a pre-trained GAN is non-isometric. The t-SNE 2D projection of $\mathcal{W}$-space of pre-trained albedo GAN confirms its non-isometric behavior as the distribution varies differently across the dimensions. Naively using the distance from the mean as a loss penalizes point $B$ more than point AA, even though $B$ is well within the distribution unlike AA. As a remedy, our proposed kNN loss uses the closeness of a point to its neighbors as a loss instead of relying on its distance from the distribution mean, promoting exploration of the entire space.
@@
where $\widebar{w}$ is the average $w$ obtained by drawing large number of random samples (typically $100k$) of $w$-codes from $\mathcal{W}$ space. This in-domain loss tries to keep the optimized latent close to the distribution mean in order to keep the estimate within the training domain. We argue that this formulation has two main drawbacks. First, using the distribution mean penalizes the solution space near the boundary of the distribution.
Second, it assumes that the distribution of $w$ is isotropic, which has no reason to hold given the use of a non-linear mapping network that maps the Gaussian distribution $\mathcal{Z}$ with the intermediate latent distribution $\mathcal{W}$ during training of StyleGAN model. We depict such limitation of in-domain loss in Fig. 3 where we first plot the distribution of $w$-codes by projecting $\mathcal{W}$-space in two dimensions using t-SNE (van der Maaten & Hinton, 2008), and validate the non-isometric behavior of $\mathcal{W}$-space. Since the distribution varies differently along different directions, one can see that simply using the distance from the mean $\widebar{w}$ as a loss (as done in case of in-domain loss) would penalize point $B$ more than point $A$ since $d_{B}>d_{A}$, even though point $B$ is well within the distribution while $A$ lies outside. Note that we use randomly sampled $100k$ $w$-codes to calculate the t-SNE projection and the distribution mean $\widebar{w}$, while the plot shows randomly selected $1000$ of them for legible visualization.
As a remedy to both the issues, we propose a novel k-Nearest Neighbors-based loss formulation that better allows the exploration of the full domain while still ensuring the estimate remains in-distribution. Instead of pushing the estimate $w_{i}$ towards the mean, we steer it closer to its neighboring $w$ codes to ensure it remains within the distribution. In non-isometric distributions like the one in Fig. 3, proximity to other samples is a more reliable indicator of a point belonging to the distribution than its distance from the mean. This approach allows the estimate to move freely within the distribution, rather than being constrained to the mean, and avoids unfairly penalizing estimates near the boundaries. We first sample a large number ($100,000$) of $z$ codes from a truncated Gaussian distribution, and pass them through the mapping network of StyleGAN2 to obtain a set $\mathcal{S_{W}}$ of $w$ codes $W$.
Figure 4: Comparison: in-domain vs. kNN loss. Left: In-domain loss attracts the estimate towards the center resulting in higher loss near the distribution boundary. Right: kNN loss ($k=5$ NN) attracts the estimate towards the nearby latents, allowing to better capture the diversity of the distribution while still keeping the estimate within the distribution. Dots represent randomly sampled latent vectors (=100) and the colors map the loss value.
@@
Then, during inversion, at each step, we obtain the $k$ nearest neighbors of the current latent code estimate $kNN(w_{i})$ from $\mathcal{S_{W}}$, and use a softmax weighted average of the distance to nearest neighbors as a loss:
$\mathcal{L}_{kNN}=\sum_{w_{i}^{k}\in kNN(w_{i})}\|w_{i}-w_{i}^{k}\|_{2}\cdot\frac{e^{-0.5\frac{\|w_{i}-w_{i}^{k}\|_{2}}{\widebar{d}^{k}_{i}}}}{D_{i}},$ (8)
@@
where $D_{i}$ is the normalizing factor, and $\widebar{d}^{k}_{i}$ is the mean distance to nearest neighbors:
$D_{i}=\sum_{w_{i}^{k}\in kNN(w_{i})}e^{-0.5\frac{\|w_{i}-w_{i}^{k}\|_{2}}{\widebar{d}^{k}_{i}}},~~~\widebar{d}^{k}_{i}=\sum_{w_{i}^{k}\in kNN(w_{i})}\|w_{i}-w_{i}^{k}\|_{2}.$ (9)
@@
Figure 4 provides an intuitive visualization of our kNN loss in 2D. Our loss landscape does not penalize as much the exploration of the distribution and its boundaries. Further, it respects the non-isometric behavior of the $\mathcal{W}$-space. The soft-max weight ensures spatial continuity and a higher contribution for neighbors closer to the optimized latent (Eq. 8). To normalize the weights, we use the mean distance to the nearest neighbors ($\bar{d_{i}^{k}}$, Eq. 9). We provide the results of our optimization scheme and ablation studies demonstrating the benefits of our kNN loss in Sec. 5. Note that we use $k=50$ for all our experiments.
@@ 4 Adapting Existing Inversion Methods
In this section we show how existing initialization and generator fine-tuning mechanisms can be adapted to JoIN to improve the inversion result and bridge the Sim-to-Real gap.
@@ 4.1 Encoder-guided Initialization
While our optimization-based approach produces faithful decompositions on synthetic data, the optimization can get stuck in local minima if the initial latent codes lie far from the target codes. To make our approach more robust, we propose to use an encoder-guided initialization inspired by hybrid approaches of GAN inversion (Chai et al., 2021b; Bau et al., 2019c; Wei et al., 2021). For each component, we can independently train a pSp network (Richardson et al., 2021a) to encode a natural image into a latent code $w$ used by the pre-trained generator to recover the given component. At test time, given a natural image, the latent code produced by the encoder can be used as an initialization for the optimization.
4.2 Generator Fine-tuning using PTI
Since our component-wise GAN models are trained using only synthetic data, they are unable to generate the features corresponding to the real images. Thus, our method poses the challenge of Sim-to-Real gap, i.e. various features of the real images are not captured faithfully in the decomposition. As a remedy, we propose to carefully fine-tune the generators in such a way that they can generate decompositions of real images while still maintaining their original priors. Indeed, approaches leveraging generator fine-tuning have become popular (Roich et al., 2021; Feng et al., 2022) where first a $w$-code is estimated using an off-the-shelf inversion technique, and then with the fixed $w$ pivot, the generator is updated to obtain close to zero reconstruction errors. Such approach is also known as Pivotal Tuning for Inversion (PTI) in the literature.
In our case, the generator fine-tuning (PTI) can be defined as:
[…]
@@ 4.2 Generator Fine-tuning using PTI
While direct fine-tuning allows the generator to overfit to the target image perfectly, it distorts the GAN manifold. In our case, allowing multiple GANs to fine-tune at once can result in loss of the priors and much greater cross-contamination (see Fig. 11(b) w/o D Loss) since we do not have any component-wise supervision for individual generators. Here, we leverage interesting fact about the presence of strong priors in GAN: since the discriminator of the GAN is trained to discriminate between real and fake samples, its knowledge about the realness of a sample can be employed to ensure the prior remains preserved during generator fine-tuning. In practice, one can use the discriminator loss to regularize the fine-tuning (Feng et al., 2022). While fine-tuning the generator around the $\widehat{w_{i}}$ pivot, we apply a new localized discriminator loss $\mathcal{L}_{LD}$, we refer to it as D Loss.
Specifically, we sample $n_{a}$ number of anchor latent codes in the close vicinity of our fixed $w$-codes by interpolating $\widehat{w_{i}}$ with randomly sampled codes $w_{k}$. At each fine-tuning step, we pass the anchor codes through the generator and compute D loss on the resulting images, keeping the discriminator frozen:
@@
where $\mathcal{L}_{G}$ is the generator loss as defined in Eq. 10. As discussed in the ablation study (Sec. 5.4), the $\mathcal{L}_{LD}$ loss ensures that images with latents close to the pivot maintain realism, preserving the generator’s priors during the optimization even as multiple generators are fine-tuned together.
@@ 5 Experiments
In this section, we present a set of experiments that showcase the efficacy of our method. We provide intrinsic image decomposition results on two different datasets: materials, and faces. We also provide results on the dataset of rendered shapes and indoor scenes in appendix. Note that for all our experiments, we use only synthetic data for training our GANs while showing results on both synthetic and real images along with comparisons with existing methods and ablation studies. We plan to release our code in future.
@@ 5.1 Preliminaries
For all our experiments we use StyleGAN2 (Karras et al., 2020) to train the generators and a pSp (Richardson et al., 2021b) encoder for initialization.
All our generators and encoders are trained using their publicly available codebases and with the standard choices of the hyperparameters.
@@
The decomposition results are evaluated using several metrics such as MSE, LPIPS, and PSNR for the synthetic images that provide ground truth. We also provide comparisons with several state-of-the-art methods that are used in the literature for intrinsic image decomposition. The results demonstrate the effectiveness of our proposed method in decomposing the input images into their intrinsic components and the ability to handle real images.
(a) Decomposition on synthetic materials.
(b) Comparisons on real materials.
Figure 5: (a) We show three examples of decomposition on synthetic materials. For each case the first column, from top to bottom, contains target image (used for optimization), groundtruth shading & albedo while the second column contains our estimates. We are able to correctly separate the highlights and to discriminate between darker shading and darker albedo. (b) Our method generalizes to real materials as well – our GAN priors allow us to better separate the highlight from the albedo leading to fewer visual artifacts. Our recovered specular is shown on the right.
@@ 5.2 Experiments on Materials Data
The Materials dataset is composed of highly realistic renderings of synthetic material tiles as done in Deschaintre et al. (2018). We obtain the albedo map directly from the material and compose the shading and specular components from blender direct and indirect rendering passes. More details on this dataset are given in the appendix (Sec. B). For this dataset, we show results for decomposition into 3 components (albedo, shading, and specular; Eq. 6) and operate at the resolution of $512\times 512$. For synthetic test images, we show results in Fig. 5(a) and quantitative comparisons to several recent SVBRDF estimation methods (Deschaintre et al., 2018; Zhou & Kalantari, 2021; Zhou & Khademi Kalantari, 2022) on the task of albedo recovery in Tab. 1. We also show competitive results on albedo map recovery both quantitatively and qualitatively on real images, taken from (Deschaintre et al., 2019), visible in Fig. 5(b). Our GAN priors allow to correctly separate flash highlights from the diffuse component and to recover plausible content under the highlights in the albedo map while correcting for the lighting fall-off. Additional results are in appendix (Sec. D).
@@ 5.3 Experiments on Faces Data
Quantitative comparisons: We obtain the reconstruction error for albedo estimation across the testset of 100 images for both the synthetic materials (in Tab. 1) and synthetic faces (in Tab. 2) datasets, and provide comparisons with existing decomposition approaches using various metrics.
Table 1: Quantitative comparisons on synthetic materials testset.
On synthetic materials Estimated Albedo
[…]
Desch18 0.021 0.314 19.53
Zhou21 0.020 0.293 20.92
Zhou22 0.021 0.296 20.18
Ours (Opt. + kNN loss) 0.014 0.326 21.58
Table 2: Quantitative comparisons on Lumos synthetic faces testset
On synthetic faces Estimated Albedo
[…]
Total Relighting 0.2220 0.0644 17.1470
PieNet 0.1843 0.0501 19.4761
InverseRenderNet++ 0.3794 0.1550 14.5227
NextFace 0.2034 0.0791 19.0699
Latent Intrinsics 0.1939 0.1053 15.7492
@@
+ kNN loss 0.1047 0.0357 23.5331
+ Encoder 0.0887 0.0345 24.1467
       + PTI w/o D Loss 0.0784 0.0155 24.1160
       + PTI w/ D Loss 0.0765 0.0072 27.3401
Figure 6: Decomposition results for Lumos faces dataset. On synthetic data, our estimated components are close to the ground truth, generator tuning is not needed when testing data is close to the GAN training distribution. Left: Results of our method for decomposing the image into two components: albedo and shading. Right: Using three components allows to recover subtle specular effects.
(a) Decomposition results on real faces.
(b) Comparisons on real face images.
Figure 7: (a) Using generator fine-tuning, our method generalizes to real faces and correctly separates albedo, shading, and specular effects while reconstructing the input correctly. (b) On real faces, our approach recovers more faithful albedo by removing shadows, and preserving the appropriate skin tones as compared to competing methods.
Faces are notoriously adapted to the usage of GANs and to cater to such a use case, we use the synthetic Lumos dataset (Yeh et al., 2022) to train our GANs, we defer more discussions on the dataset to the appendix.
First, we show results on test images from the dataset in Fig. 6(left) and Fig. 6(right) where we use respectively two and three components. For this experiment, we use the encoder initialization and our optimization without generator fine-tuning (PTI). As can be seen in these figures our method is able to recover accurate albedo, shading, and specular maps and does not suffer from cross-channel contamination.
As the synthetic dataset does not reproduce the complexity and diversity of real faces, for real images, we add the generator fine-tuning (w/ local D loss) step presented in Sec. 4.2. We show results on real faces taken from the FFHQ dataset in Fig. 7(a) and compare our method to five recent relighting methods, Total Relighting (Pandey et al., 2021), Lumos (Yeh et al., 2022), InverseRenderNet++ (Yu & Smith, 2021), PieNet (Das et al., 2022a), Nextface (Dib et al., 2021a; b; 2022), and Latent Intrinsics (Zhang et al., 2024) that seek to recover face albedo as part of their relighting pipeline, as shown in Fig. 7(b). In all the cases, the predicted components by our method maintain the priors of each GAN. We can observe that our method works better at producing shading-free albedo images that exhibit fewer shading variations around the neck, nose, nostrils, and mouth.
For quantitative comparisons, we use a testset of synthetic images from Lumos dataset, and compare reconstruction errors for albedo recovery for all variants of our approach with four competing methods in Tab. 2. We see that our method achieves the lowest errors. Additional results are provided in appendix (Sec. D).
Application to relighting. GANs are known for their rich image manipulation capabilities. Our modular approach allows leveraging such capabilities to modify each image component separately using GAN’s latent space editing, making it suitable for image modifications such as relighting. We show real image relighting results in Fig. 8. To generate these results, after obtaining the inversions for albedo, shading, and specular, we use an off-the-shelf latent space editing method SeFa (Shen & Zhou, 2021) on the shading generator to modify the shading image while keeping the other two components fixed. StyleGAN’s disentangled representation helps with identity preservation. As seen in Fig. 8, our method produces varied novel lighting.
Figure 8: Face relighting results. Once the decomposition of a real face is obtained, we modify the shading component using GAN editing technique Shen & Zhou (2021) on the shading generator obtaining plausible relighting edits.
Table 3: Effectiveness of kNN Loss
in inverting a single GAN
On CelebA-HQ test set Inversion on single GAN trained on FFHQ
[…]
Optimization $0.384$ $0.095$ $16.908$
Opt. + in-domain loss $0.384$ $0.094$ $16.990$
Opt. + kNN loss (ours) $\mathbf{0.379}$ $\mathbf{0.092}$ $\mathbf{17.120}$
Table 4: Quantitative comparisons for different variations of our approach on synthetic Faces.
On synthetic Faces (Lumos dataset)
Albedo Shading Specular Image
[…]
Inversion w/ single GAN for both albedo and shading 0.1954 0.1269 14.166 0.2657 0.07021 18.245 - - - 0.09342 0.03133 24.466
Optimization 0.1311 0.0417 22.476 0.2832 0.0518 18.863 0.2544 0.0336 32.320 0.0879 0.0308 24.296
Opt. + in-domain loss 0.1209 0.0400 22.589 0.2738 0.0497 19.159 0.2444 0.0332 32.833 0.0876 0.0302 24.164
Opt. + kNN loss 0.1047 0.0357 23.533 0.2603 0.0487 20.911 0.2361 0.0320 32.905 0.0868 0.0301 24.114
Encoder + Opt. + kNN loss 0.0887 0.0345 24.146 0.2039 0.0414 21.032 0.1722 0.0282 33.409 0.0713 0.0306 26.445
Encoder + Opt. + kNN loss + PTI w/o D Loss 0.0784 0.0155 24.116 0.2775 0.0519 18.165 0.1550 0.0278 32.838 0.0136 0.0011 33.545
Encoder + Opt. + kNN loss + PTI w/ D Loss 0.0765 0.0072 27.340 0.1703 0.0285 22.475 0.0774 0.0251 34.547 0.0194 0.0022 32.517